Maria has opened her own carpentry business. She gave up a job earning $3000 a month to do this. She has to pay $2000 per month on rent for her shop. She also purchased tools and equipment that she took a loan for and she now pays $500 a month for the loan. This month Maria has also spent $1000 on woo, nails, and other variable costs.

7. What is Maria’s implicit cost this month?

8. What will be her total fixed costs this month?

9. What are her total explicit costs this month?

10. If Maria sells $8,000 worth of goods (revenue) what are her accounting profits for the month?

11. With revenue of $8,000, what are her economic profits for the month?
Answer #1

Explicit cost = Direct cost

Explicit cost = 2000 + 1000 + 500= $ 3500

Implicit cost = Indirect or, opportunity cost

Implicit cost = $ 3,000

7. Implicit cost = $ 3,000

8. Total fixed cost = $ 500

She has taken a loan against which maria has to par $ 500 per month. This has yo be paid even though she does not produce and sell anything. Therefore, this is fixed cost.

9. Explicit cost = $ 3,500

10. Accounting profit = Revenue - Explicit cost

=> Accounting profit = $ 8,000 - $ 3,500= $ 4,500

11. Economic profit = Accounting profit - Implicit cost

=> Economic profit = $ 4,500 - 3,000 = $ 1500